# See What You Seek: Semantic Contextual Integration for Cloth-Changing Person Re-Identification

### v1

@@ IV-B Implementation Details
We adopt a modified ResNet-50 pre-trained on CLIP as the backbone for feature extraction. Images are resized to pixels, with a batch size of 64. Training employs the Adam optimizer [44] with random horizontal flipping, cropping, and erasing [45] for data augmentation. A global attention pooling layer reduces the feature dimension from 2048 to 1024, matching the text feature dimension, which is scaled from 512 to 1024.
In the first stage, two prompt learners are trained for 60 epochs on LTCC, VC-Clothes, Market-1501, and MSMT17, and 20 epochs on PRCC, with an initial learning rate of adjusted using a cosine decay schedule. In the second stage, the visual encoder is trained for 120 epochs on LTCC, VC-Clothes, Market-1501, and MSMT17, and 20 epochs on PRCC, with learning rates initially set to and reduced by a factor of 10 at the 40th and 70th epochs. All experiments are conducted on the PyTorch platform using a single A100 GPU.
@@ IV-C Comparison with State-of-the-Art Methods
We compare our method with state-of-the-art techniques on LTCC, PRCC, and VC-Clothes, including traditional Re-ID methods HACNN [25], PCB [24], IANet [26], and ISP [27], and specialized CC-ReID methods FSAM [33], RCSANet [35], CAL [10], GI-ReID [12], M2Net [34], and AIM [11].
@@
| SCI (Ours) | – | 75.7 | 40.6 | 42.1 | 18.6 | 99.6 | 97.7 | 59.8 | 56.2 | 94.9 | 89.2 | 89.2 | 83.1 |
As shown in Table I, CC-ReID methods significantly outperform general Re-ID methods, primarily because general methods rely heavily on clothing for identification, which is less effective in long-term surveillance with clothing changes. Our SCI method achieves competitive results by effectively eliminating ID-irrelevant features and enhancing positive ID elements in visual images.
On LTCC, SCI surpasses the baseline [17] in the cloth-changing setting, with improvements of 3.8% in Rank-1 accuracy and 2.7% in mAP. Although our mAP is slightly lower than AIM [11] in the general setting, this is likely due to LTCC’s inherent challenges, where models often rely on clothing information. By filtering out such information, SCI experiences minor declines in non-cloth-changing scenarios, but these are negligible given the overall benefits. Notably, SCI achieves these results without incorporating additional modalities or modules, unlike other methods. The primary advantage of our method lies in its ability to avoid carrying clothing-relevant information, which is often unavoidable in methods relying on auxiliary data such as poses and gaits. Compared with non-auxiliary-based methods, SCI outperforms CAL by 12.7% in Rank-1 and 6.3% in mAP in the cloth-changing setting on VC-Clothes. On PRCC, SCI also surpasses AIM by 3.5% in Rank-1 accuracy for clothing changes. Although the differences on other metrics are less pronounced due to PRCC’s simpler structure (only two outfits per ID), SCI still shows notable improvements. On VC-Clothes, SCI demonstrates significant gains in the cloth-changing setting, with a 7.4% higher mAP than AIM, underscoring the effectiveness of our approach in challenging scenarios.
@@ IV-D Ablation Studies and Analysis
We conduct ablation studies to assess the effectiveness of the components in SCI and analyze parameter impacts.
IV-D1 Effectiveness of Semantic Separation Enhancement
We first evaluate the baseline without additional modules, where the model relies solely on a prompt “a photo of a person” for classification. This approach shows reduced performance. Introducing the SSE module, which uses two prompts (“a photo of a person” and “a photo of the clothes”) to learn features separately at the textual level, improves Rank-1 and mAP by around 2.0% across datasets, demonstrating its ability to filter clothing features.
@@ IV-D1 Effectiveness of Semantic Separation Enhancement
| Baseline | SSE | SIM | LTCC | PRCC | VC-Clothes | |||||||||
| General | Cloth-changing | Same-clothes | Cloth-changing | General | Cloth-changing | |||||||||
| R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | |||
| ● | ○ | ○ | 73.0 | 36.5 | 38.3 | 15.9 | 100.0 | 99.5 | 57.0 | 55.3 | 93.1 | 86.9 | 85.9 | 79.3 |
| ● | ● | ○ | 76.5 | 39.7 | 39.8 | 18.3 | 100.0 | 98.4 | 59.0 | 54.7 | 94.1 | 88.0 | 87.6 | 81.1 |
| ● | ○ | ● | 72.8 | 36.3 | 37.8 | 16.1 | 99.7 | 98.0 | 52.4 | 49.9 | 94.0 | 88.9 | 88.4 | 82.4 |
| ● | ● | ● | 75.7 | 40.6 | 42.1 | 18.6 | 99.6 | 97.7 | 59.8 | 56.2 | 94.9 | 89.2 | 89.2 | 83.1 |
IV-D2 Effectiveness of Semantic-Guided Interaction Module
We further assess SIM by integrating it with the baseline model. As shown in Table II, using SIM alone often degrades performance, particularly in the cloth-changing scenario on PRCC, where Rank-1 drops by 4.6% and mAP decreases from 55.3% to 49.9%.
@@ IV-D3 Feature Distribution Analysis
From Figs. 4(a) and (c), it is observable that both the baseline and our method exhibit disorganized feature distributions with blurred class boundaries. During this stage, only the prompts are trained while the visual and text encoders remain frozen, limiting their capability to effectively extract diverse representations. In contrast, as shown in Figs. 4(b) and (d), the feature distributions for both methods become clearer, indicating the effectiveness of the representations derived from the text-image similarity computations.
@@
IV-D4 Visualization of Feature Maps
To illustrate the distinctions between the baseline [17] and SCI, we present heatmaps of regions of interest in Fig. 5, demonstrating performance across LTCC, VC-Clothes, and PRCC. Fig. 6 shows VC-Clothes for the same ID but with different outfits, illustrating the regions of interest that our SCI model focuses on under clothing variations. These visualizations provide insights into the unique differences in feature activation patterns.
@@ IV-D4 Visualization of Feature Maps
| (a) R-1 Accuracy on cloth-changing | (b) mAP Accuracy on cloth-changing | (c) R-1 Accuracy on general | (d) mAP Accuracy on general |
IV-D5 Comparative Analysis of Mechanisms in SIM
As shown in Fig. 7, we evaluate the effectiveness of the non-local operation [41] compared to self-attention mechanisms [40] in leveraging textual information obtained from the SSE module to guide visual representations. In cloth-changing scenarios, the non-local mechanism yields superior performance, achieving 42.1% in Rank-1 accuracy and 18.6% in mAP, as depicted in Figs. 7(a) and (b), thereby outperforming the self-attention mechanism by margins of 4.6% and 2.4%, respectively.
This underscores the non-local mechanism’s enhanced capability to capture comprehensive global context, which is vital for integrating information from disparate yet significant segments of the visual field in cloth-changing scenarios. In general scenarios, as illustrated in Figs. 7(c) and (d), both mechanisms exhibit comparable efficacy, with the non-local mechanism slightly edging out self-attention by a 2.0% increase in mAP. This suggests the non-local mechanism is robust and can be effectively deployed for person Re-ID tasks.
IV-D6 Loss Functions Analysis
We conduct experimental evaluations to explore the effectiveness of (9), as demonstrated in Fig. 8(a). This loss function significantly improves performance by optimizing for higher ID similarity and reducing clothing similarity, making it particularly effective for applications that need to distinguish personal ID from external features like clothing, relevant in tasks such as cloth-changing scenarios. It helps the model focus on positive factors while ignoring negative variations such as clothing changes, thereby enhancing adaptability and accuracy in complex environments.
@@ IV-D6 Loss Functions Analysis
| SCI (Ours) | – | 85.7 | 68.6 | 73.5 | 47.4 | |
IV-D7 Cloth-consistent Re-ID Scenario
To validate the applicability of our method in cloth-consistent scenarios, we conducted experiments using our cloth-changing approach on two traditional ReID datasets: Market1501 and MSMT17. For ease of comparison, we also incorporated cloth-consistent methods: AGW [48], reID-NAS [49], CLIP-ReID [17], and IRM [50]. This enables a comprehensive evaluation of the performance in varying scenarios.
@@ IV-D7 Cloth-consistent Re-ID Scenario
As shown in Table III (cloth-consistent), existing Re-ID methods designed for traditional datasets have achieved progressively better performance, even on larger-scale datasets like MSMT17. While our method is specifically designed to address the more realistic scenario of clothing changes, we are also interested in evaluating its performance on conventional datasets. From Table III (cloth-changing), it can be observed that our SCI method achieves the best performance on both Market1501 and MSMT17 among cloth-changing methods. Although there is still a gap compared to cloth-consistent methods, this highlights an important direction for future research and improvement.
IV-D8 Visualization of Retrieved Examples
We present the visualization of the top-10 retrieved results from randomly selected query examples on LTCC in Fig. 9. It observes that the retrieval results are predominantly optimistic, reflecting a high level of accuracy and relevance in most cases. However, some incorrect results still persist, which could be due to challenges such as varying viewpoints, low lighting conditions, or occlusions within the images. These issues underscore potential areas for future research. Despite these challenges, the overall result strongly supports the effectiveness of our SCI, demonstrating its robustness across a variety of scenarios.
IV-D9 Parameter Selection
@@
We evaluate the parameters and specified in (9). The results for LTCC are displayed in Fig. 8(b). We determine that and provide the best performance in most scenarios, optimizing the balance between enhancing ID and reducing clothing similarity, while feature orthogonalization minimizes interference between ID and clothing features. This setting allows the model to effectively discern between the positive representations and their negative factors, such as clothing. Additionally, and obtain the second-best results. This analysis underscores the effectiveness and adaptability of our proposed method in real-world surveillance applications.
IV-D10 Efficiency Analysis
@@
To verify the efficiency and performance of SCI, we analyze several computational and training metrics across different methods, detailed in Table IV. AIM [11] utilizes a dual-branch ResNet model, achieving an R-1 accuracy of 41.8%, which surpasses CAL [10], another ResNet-based model. Although AIM requires higher GFLOPs compared to CAL, it significantly outperforms CAL in terms of accuracy with the same parameter count.
In CLIP-ReID [17] and our method, we integrate textual modality information, transitioning from a single-modality to a multimodal approach. This justifies the observed increases in both GFLOPs and parameters, which remain within acceptable limits. Our SCI achieves an R-1 accuracy of 42.1%, demonstrating the value of incorporating textual information. This allows the model to better distinguish between different IDs, especially in challenging scenarios. Overall, our method provides a good balance between computational cost and performance enhancements.

### v2

@@ IV-B Implementation Details
We adopt two backbones for feature extraction: a modified ResNet-50 and ViT-B/16 pre-trained on CLIP. All images are resized to pixels, with a batch size of 64. We use the Adam optimizer [62] and apply random horizontal flipping, cropping, and erasing [63] for data augmentation.
For the CNN branch, a global attention pooling layer to reduce the feature dimension from 2048 to 1024, matching the text feature dimension (scaled from 512 to 1024). For the transformer branch, we employ 12 transformer layers with hidden size 768; a linear projection reduces the image feature dimension from 768 to 512 to align with the text features.
Training proceeds in two stages: 1) Stage 1 (prompt learning): train two prompt learners for 60 epochs (ResNet-50) or 120 epochs (ViT-B/16) with an initial learning rate of and cosine decay. 2) Stage 2 (visual encoder fine-tuning): train the visual encoder for 120 epochs (ResNet-50) or 30 epochs (ViT-B/16), starting at and reducing by a factor of 10 at epochs 40 and 70.
All experiments are implemented in PyTorch and run on a single NVIDIA A100 GPU.
@@ IV-C Comparison with State-of-the-Art Methods
We compare SCI against state-of-the-art methods on LTCC, PRCC and VC-Clothes, including general ReID approaches (ISP [24], CLIP-ReID [14], and Instruct-ReID [15]) and specialized cloth-changing methods (FSAM [32], RCSANet [44], CAL [7], GI-ReID [9], M2Net [33], AIM [8], Pos-Neg [42], ACID [64], 3DInvarReID [39], DCR-ReID [37], IMS+GEP [46], AFL [4], CVSL [36], MCSC [45], MAL-F [41], and DLCR [47]).
@@
Instruct-ReID [15] | CVPR’24 | 75.8 | 52.0 | - | - | - | - | 54.2 | 52.3 | - | - | 89.7 | 78.9 |
AFL [4] | TMM’24 | 74.4 | 39.1 | 42.1 | 18.4 | 100.0 | 99.7 | 57.4 | 56.5 | 93.9 | 88.3 | 82.5 | 83.0 |
MCSC [45] | TIP’24 | 73.9 | 40.2 | 42.2 | 19.4 | 99.8 | 99.8 | 57.8 | 57.3 | 93.2 | 87.9 | 83.3 | 83.2 |
IMS+GEP [46] | TMM’24 | - | - | 43.4 | 18.2 | 99.7 | 99.8 | 57.3 | 65.8 | 94.7 | 94.9 | 81.8 | 81.7 |
CVSL [36] | WACV’24 | 76.4 | 41.9 | 44.5 | 21.3 | 97.5 | 99.1 | 57.5 | 56.9 | - | - | - | - |
MAL-F [41] | TITS’25 | 37.5 | 16.2 | 38.8 | 17.0 | - | - | 57.1 | 57.7 | - | - | - | - |
DLCR [47] | WACV’25 | - | - | 41.3 | 19.6 | - | - | 66.5 | 63.0 | - | - | 87.1 | 81.1 |
SCI w/ ResNet-50 | Ours | 75.7 | 40.6 | 42.1 | 18.6 | 99.6 | 97.7 | 59.8 | 56.2 | 94.9 | 89.2 | 89.2 | 83.1 |
SCI w/ ViT-B | Ours | 79.9 | 44.4 | 45.9 | 21.6 | 100.0 | 99.4 | 62.8 | 60.0 | 95.4 | 91.6 | 91.7 | 87.5 |
As shown in Table I, SCI consistently outperforms baselines by effectively suppressing ID-irrelevant features and reinforcing discriminative ID cues.
IV-C1 CNN-Based Model
On LTCC (cloth-changing), SCI improves Rank-1 by 3.8% and mAP by 2.7% over CLIP-ReID [14]. Although SCI’s mAP is marginally lower than AIM [8] in the general setting, this is expected: by filtering out clothing information, SCI may sacrifice some performance when clothing cues are reliable, but gains robustness under cloth variation.
On PRCC, SCI surpasses AIM by 3.5% in Rank-1 accuracy under clothing changes. Due to PRCC’s simpler two-outfit per ID setup, improvements on other metrics are relatively smaller but still notable.
On VC-Clothes, SCI achieves a substantial mAP gain of 7.4% over AIM in the cloth-changing setting, further validating its effectiveness in challenging scenarios.
IV-C2 ViT-Based Model
With ViT-B/16, SCI attains near state-of-the-art performance, especially on LTCC, reaching 45.9% Rank-1 under severe clothing variations. Crucially, these results are obtained without auxiliary models or additional preprocessing, distinguishing SCI from methods that rely on external components.
@@ IV-D Ablation Studies and Analysis
We perform ablation studies to assess each module’s contribution and analyze key parameters.
Backbone | Baseline | SSE | SIM | LTCC | PRCC | VC-Clothes | |||||||||
General | Cloth-changing | Same-clothes | Cloth-changing | General | Cloth-changing | ||||||||||
R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | ||||
ResNet-50 | ● | ○ | ○ | 73.0 | 36.5 | 38.3 | 15.9 | 100.0 | 99.5 | 57.0 | 55.3 | 93.1 | 86.9 | 85.9 | 79.3 |
● | ● | ○ | 76.5 | 39.7 | 39.8 | 18.3 | 100.0 | 98.4 | 59.0 | 54.7 | 94.1 | 88.0 | 87.6 | 81.1 | |
● | ○ | ● | 72.8 | 36.3 | 37.8 | 16.1 | 99.7 | 98.0 | 52.4 | 49.9 | 94.0 | 88.9 | 88.4 | 82.4 | |
● | ● | ● | 75.7 | 40.6 | 42.1 | 18.6 | 99.6 | 97.7 | 59.8 | 56.2 | 94.9 | 89.2 | 89.2 | 83.1 | |
ViT-B | ● | ○ | ○ | 79.5 | 43.6 | 41.6 | 19.4 | 99.9 | 99.2 | 60.9 | 58.5 | 94.4 | 89.8 | 90.6 | 84.7 |
● | ● | ○ | 76.3 | 43.9 | 43.4 | 21.0 | 100.0 | 99.3 | 62.3 | 58.7 | 94.6 | 89.8 | 91.4 | 84.8 | |
● | ○ | ● | 77.5 | 43.9 | 42.9 | 20.5 | 99.8 | 99.2 | 61.4 | 59.1 | 94.3 | 90.4 | 90.9 | 85.7 | |
● | ● | ● | 79.9 | 44.4 | 45.9 | 21.6 | 100.0 | 99.4 | 62.8 | 60.0 | 95.4 | 91.6 | 91.7 | 87.5 |
IV-D1 Impact of Semantic Separation Enhancement
We first evaluate the baseline without additional modules, where the model relies solely on a prompt “a photo of a person” for classification. This approach shows reduced performance. Introducing the SSE module, which uses two prompts (“a photo of a person” and “a photo of the clothes”) to learn features separately at the textual level. We observe consistent improvements of approximately 2.0% in both Rank-1 accuracy and mAP across datasets for both CNN-based and ViT-based models, as shown in Table II. This demonstrates the effectiveness of the SSE module in filtering clothing features.
@@ IV-D1 Impact of Semantic Separation Enhancement
IV-D2 Impact of Semantic-Guided Interaction Module
We further assess SIM by integrating it with the baseline model. As shown in Table II, using SIM alone often degrades performance based on CNN, particularly in the cloth-changing scenario on PRCC, where Rank-1 drops by 4.6% and mAP decreases from 55.3% to 49.9%.
@@ IV-D3 Feature Distribution via t-SNE
From Fig. 4(a) and (c), it is observable that both the baseline and our method exhibit disorganized feature distributions with blurred class boundaries. During this stage, only the prompts are trained while the visual and text encoders remain frozen, limiting their capability to effectively extract diverse representations. In contrast, as shown in Fig. 4(b) and (d), the feature distributions for both methods become clearer, indicating the effectiveness of the representations derived from the text-image similarity computations.
@@
IV-D4 Qualitative Prompt-Guided Refinement
To illustrate the distinctions between the baseline [14] and SCI, we present in Fig. 5 the attention maps activated by different prompts, highlighting how prompts guide the refinement of visual features across LTCC, VC-Clothes, and PRCC. It demonstrates that the learned prompts are capable of guiding the network to focus on relevant semantic areas.
Fig. 6 shows VC-Clothes for the same ID but with different outfits, illustrating the regions of interest that our SCI model focuses on under clothing variations. These visualizations provide insights into the unique differences in feature activation patterns. Moreover, while SCI leverages prompt learning to jointly model ID and clothing information, it fundamentally differs from attribute-based approaches, as it neither explicitly recognizes attributes nor relies on additional attribute detectors.
@@ IV-D4 Qualitative Prompt-Guided Refinement
(a) distribution. | (b) distribution. | (c) distribution. |
IV-D5 Visualization of Feature Similarity
In Fig. 7, we visualize the similarity distributions of features between the text prompts and the images. Fig. 7(a) is derived from the prompt in Eq. 1, while the distribution is relatively well-formed, intra-class similarity remains dispersed. Fig. 7(b) shows the distribution from the prompt in Eq. 2, which learns clothing-related semantics. However, due to significant clothing variations over time for the same individual, intra-class differences outweigh inter-class differences. Fig. 7(c) shows the distribution after orthogonalizing the features, which achieves a clearer separation between intra- and inter-class similarity, indicating improved feature discrimination. This demonstrates the effectiveness of dual textual prompts and orthogonalization in enhancing feature alignment.
(a) R-1 on cloth-changing | (b) mAP on cloth-changing | (c) R-1 on general | (d) mAP on general |
IV-D6 Comparative Analysis of Mechanisms in SIM
As shown in Fig. 8, we evaluate the effectiveness of the non-local operation [59] compared to self-attention mechanisms [58] in leveraging textual information obtained from the SSE module to guide visual representations. In cloth-changing scenarios, the non-local mechanism yields superior performance, achieving 42.1% in Rank-1 accuracy and 18.6% in mAP, as depicted in Fig. 8(a) and (b), thereby outperforming the self-attention mechanism by margins of 4.6% and 2.4%, respectively.
This underscores the non-local mechanism’s enhanced capability to capture comprehensive global context, which is vital for integrating information from disparate yet significant segments of the visual field in cloth-changing scenarios. In general scenarios, as illustrated in Fig. 8(c) and (d), both mechanisms exhibit comparable efficacy, with the non-local mechanism slightly edging out self-attention by a 2.0% increase in mAP. This suggests the non-local mechanism is robust and can be effectively deployed for person ReID tasks.
IV-D7 Parameter Sensitivity
We conduct experimental evaluations to explore the effectiveness of Eq. 8, as demonstrated in Fig. 9(a). This loss function significantly improves performance by optimizing for higher ID similarity and reducing clothing similarity, making it particularly effective for applications that need to distinguish personal ID from external features like clothing, relevant in tasks such as cloth-changing scenarios. It helps the model focus on positive factors while ignoring negative variations such as clothing changes, thereby enhancing adaptability and accuracy in complex environments.
IV-D8 Generalization to Cloth-Consistent Scenarios
To validate the applicability of our method in cloth-consistent scenarios, we conducted experiments using our cloth-changing approach on two traditional ReID datasets: Market1501 and MSMT17. For ease of comparison, we also incorporated cloth-consistent methods: AGW [22], TransReID [23], CLIP-ReID [14], and Instruct-ReID [15]. This enables a comprehensive evaluation of the performance in varying scenarios.
@@ IV-D8 Generalization to Cloth-Consistent Scenarios
3DInvarReID [39] | ICCV’23 | 95.1 | 97.9 | 80.8 | 59.1 | |
AFL [4] | TMM’24 | 95.5 | 88.8 | 80.9 | 58.1 | |
FIRe2 [40] | TIFS’24 | 95.4 | 87.7 | 79.7 | 56.2 | |
SCI w/ ResNet-50 | Ours | 94.4 | 86.5 | 81.4 | 59.5 | |
SCI w/ ViT-B | Ours | 95.2 | 89.4 | 89.6 | 74.7 |
Methods | Venue | Input Resolution | Params (M) | GFLOPs | Training Time (min) | FPS | R-1 | ||
LTCC | PRCC | VC-Clothes | |||||||
CAL [7] | CVPR’22 | 23.5 | 9.2 | 41 | 538 | 38.0 | 54.7 | 82.6 | |
CLIP-ReID [14] | AAAI’23 | 77.6 | 10.6 | 67 | 314 | 38.3 | 57.0 | 85.9 | |
3DInvarReID [39] | ICCV’23 | 217.6 | – | 241 | – | 40.9 | 56.5 | - | |
AIM [8] | CVPR’23 | 72.7 | 9.2 | 52 | 754 | 41.8 | 56.3 | 81.0 | |
SCI w/ ResNet-50 | Ours | 93.9 | 10.6 | 78 | 302 | 42.1 | 59.8 | 89.2 | |
SCI w/ ViT-B | Ours | 143.7 | 26.7 | 46 | 80 | 45.9 | 62.8 | 91.7 |
@@
As shown in Table III (cloth-consistent), existing ReID methods designed for traditional datasets have achieved progressively better performance, even on larger-scale datasets like MSMT17. Although our proposed SCI framework is primarily designed to address the more practical scenario of clothing changes, we also evaluate its effectiveness on conventional datasets to assess its generalization ability. From Table III (cloth-changing), it can be observed that our SCI method achieves the best performance on both Market1501 and MSMT17 among cloth-changing methods. In particular, the ViT-based SCI achieves a Rank-1 accuracy of 89.6% on MSMT17, outperforming Instruct-ReID by 2.7% in Rank-1 and 2.3% in mAP. These results underscore the strong generalization capability of our approach.
IV-D9 Visualization of Retrieved Examples
We present the visualization of the top-10 retrieved results from randomly selected query examples on LTCC in Fig. 10. It observes that the retrieval results are predominantly optimistic, reflecting a high level of accuracy and relevance in most cases.
However, some incorrect results still persist, as shown in Fig. 11, mainly due to limited views and highly similar clothing colors, which make retrieval harder. This problem is common when people wear uniforms or dark winter clothes. Further improvements are needed to enhance the model’s ability to distinguish individuals with similar appearances.
IV-D10 Parameter Selection
@@ IV-D9 Visualization of Retrieved Examples
We evaluate the parameters and specified in Eq. 8. The results for LTCC are displayed in Fig. 9(b). We determine that and provide the best performance in most scenarios, optimizing the balance between enhancing ID and reducing clothing similarity, while feature orthogonalization minimizes interference between ID and clothing features. This setting allows the model to effectively discern between the positive representations and their negative factors, such as clothing. Additionally, and obtain the second-best results. This analysis underscores the effectiveness and adaptability of our proposed method in real-world surveillance applications.
IV-D11 Efficiency Comparison
Table IV presents a comprehensive comparison of parameter size, computational complexity (GFLOPs), training time, inference speed (FPS), and Rank-1 accuracy across representative ReID methods.
@@
AIM [8] utilizes a dual-branch ResNet model, achieving an Rank-1 accuracy of 41.8%, which surpasses CAL [7], another CNN-based model. Although AIM requires higher parameters compared to CAL, it significantly outperforms CAL in terms of accuracy with the same parameter count.
In CLIP-ReID [14] and our method, we integrate textual modality information, transitioning from a single-modality to a multimodal approach. This justifies the observed increases in both GFLOPs and parameters, which remain within acceptable limits. Our SCI w/ ResNet-50 incorporates a prompt learning strategy and achieves 42.1% Rank-1 on LTCC, surpassing CLIP-ReID by 3.8%, while maintaining similar computational efficiency (10.6 GFLOPs vs. 10.6, and 302 FPS vs. 314). This highlights the effectiveness of our design in utilizing textual guidance without introducing significant overhead. Furthermore, our SCI w/ ViT-B achieves the best performance across all datasets, achieving 45.9% Rank-1 accuracy on LTCC, 62.8% on PRCC, and 91.7% on VC-Clothes. Although it has higher computational requirements due to the ViT backbone (143.7M parameters, 26.7 GFLOPs), the performance gains are substantial. Overall, SCI achieves a good trade-off between accuracy and efficiency, making it well-suited for practical deployment in real-world scenarios.